OpenStudy (anonymous):

Which shows three ordered pairs that are all solutions of the equation y = x + 7? A. (0, 7), (2, 9), and (3, 11) B. (0, 7), (2, 9), and (3, 10) C. (1, 7), (2, 9), and (3, 10) D. (1, 8), (2, 9), and (3, 11)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@alekos

hartnn (hartnn):

plug in the points in the equation like : 0,7 y=x+7 7= 0+7 7=7 true so 0,7 satisfies the equation and hence lies on y=x+7 do it ofr other points too